OpenStudy (anonymous):

let g(X) = abs(cosx-1) the maximum value attained by g on the closed interval [0, 2pi] is for x equal to a. -1 b. 0 c. pi/2 d. 2 e. pi

OpenStudy (anonymous):

It is 2. for x=pi... cos x -1 = -2... so abs (Cos x -2) = 2

OpenStudy (anonymous):

d. 2

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

@muah 143 didn't you see my explanation for the answer?

OpenStudy (anonymous):

yeah i did...i was actually trying to see RyanL. point of view

OpenStudy (anonymous):

for any real number \[-1\le\cos x \le1\]\[-2\le\cos x -1\le0\]\[0\le|\cos x -1|\le2\]

OpenStudy (anonymous):

ohhh thanks

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

:)